# Nonreciprocal transmission in hybrid atomic ensemble-optomechanical systems

### v3

@@ III.3 Conditions for perfect nonreciprocity
Figure 5: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of $y/\kappa_{2}$ for different values of atomic ensemble decay rate $f$. (a) $f/\kappa_{2}=0.1$, (b) $f/\kappa_{2}=1$, (c) $f/\kappa_{2}=5$. The other parameters are the same as in Figure 2, except for $\gamma/\kappa_{2}=0.01$, $\kappa_{1}/\kappa_{2}=10$, and $J_{3}=\sqrt{\frac{-R_{8}+\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$.
Figure 6: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of $y/\kappa_{2}$ for different values of atomic ensemble decay rate $f$. (a) $f/\kappa_{2}=0.1$, (b) $f/\kappa_{2}=1$, (c) $f/\kappa_{2}=5$. The other parameters are the same as in Figure 2, except for $\gamma/\kappa_{2}=0.01$, $\kappa_{1}/\kappa_{2}=10$, and $J_{3}=-\sqrt{\frac{-R_{8}-\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$.
@@ III.4 Influence of atomic ensemble decay and mechanical damping
Figure 7: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning y$/\kappa_{2}$ for different values of mechanical damping rate $\gamma$. (a) $\gamma/\kappa_{2}=0.001$, (b) $\gamma/\kappa_{2}=0.01$, (c) $\gamma/\kappa_{2}=0.1$, (d) $\gamma/\kappa_{2}=1$. The other parameters are the same as in Figure 2, except for $f/\kappa_{2}=1$, $J_{3}=\sqrt{\frac{-R_{8}+\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$ and $\kappa_{1}/\kappa_{2}=10$.
Figure 8: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning y$/\kappa_{2}$ for different values of mechanical damping rate $\gamma$. (a) $\gamma/\kappa_{2}=0.001$, (b) $\gamma/\kappa_{2}=0.01$, (c) $\gamma/\kappa_{2}=0.1$, (d) $\gamma/\kappa_{2}=1$. The other parameters are the same as in Figure 2 except for $f/\kappa_{2}=1$, $J_{3}=-\sqrt{\frac{-R_{8}-\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$ and $\kappa_{1}/\kappa_{2}=10$.
@@ Acknowledgments
$\mathcal{D}$ $=2(-2J_{1}|J_{2}|J_{3}G_{2}\cos(\theta-\phi)-2i|J_{2}|J_{3}\kappa_{2}G_{1}\cos\phi$ (31)
@@
$+y^{2}\mathcal{D}_{7}+\kappa_{1}\kappa_{2}\mathcal{D}_{8}-i\gamma fy\mathcal{D}_{9}+y^{4}),$
@@
For simplicity, we consider the case where $\theta=\phi=\frac{\pi}{2}$, and the expression of $\mathcal{D}$ becomes $\mathcal{D}=2(J_{3}^{2}(R_{5}+R_{2}^{\prime})+R_{4}+\frac{R_{6}}{J_{3}^{2}}(J_{1}J_{3}^{2}+J_{1}\gamma f-G_{1}G_{2}f)^{2}+R_{3})$, where $R_{1}=\frac{1}{\sqrt{\kappa_{1}\kappa_{2}}}$, $R_{2}=\frac{-2G_{1}G_{2}f}{R_{1}}$, $R^{\prime}_{2}=J_{1}^{2}+\kappa_{1}\kappa_{2}$, $R_{3}=G^{2}_{2}f\kappa_{1}+G_{1}^{2}f\kappa_{2}+J_{1}^{2}\gamma f+\kappa_{1}\kappa_{2}\gamma f$, $R_{4}=-2J_{1}f(J_{1}\gamma-G_{1}G_{2})$, $R_{5}=-2J_{1}^{2}$, $R_{6}=\frac{G_{2}^{2}+\gamma\kappa_{2}}{G_{2}^{2}}$, $R_{7}=R_{5}+R_{2}^{\prime}+R_{6}J_{1}^{2}$, $R_{8}=R_{4}-R_{2}+2R_{6}J_{1}^{2}\gamma f-2R_{6}J_{1}G_{1}G_{2}f+R_{3}$, and $R_{9}=R_{6}J_{1}^{2}\gamma^{2}f^{2}-2R_{6}J_{1}\gamma f^{2}G_{1}G_{2}+R_{6}G_{1}^{2}G_{2}^{2}f^{2}$.

### v4

@@ III.3 Conditions for perfect nonreciprocity
Figure 5: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of $y/\kappa_{2}$ for different values of atomic ensemble decay rate $f$. (a) $f/\kappa_{2}=0.1$, (b) $f/\kappa_{2}=1$, (c) $f/\kappa_{2}=5$. The other parameters are the same as in Figure 2, except for $\gamma/\kappa_{2}=0.01$, $\kappa_{1}/\kappa_{2}=1$, and $J_{3}=\sqrt{\frac{-R_{8}+\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$.
Figure 6: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of $y/\kappa_{2}$ for different values of atomic ensemble decay rate $f$. (a) $f/\kappa_{2}=0.1$, (b) $f/\kappa_{2}=1$, (c) $f/\kappa_{2}=5$. The other parameters are the same as in Figure 2, except for $\gamma/\kappa_{2}=0.01$, $\kappa_{1}/\kappa_{2}=1$, and $J_{3}=-\sqrt{\frac{-R_{8}-\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$.
@@ III.4 Influence of atomic ensemble decay and mechanical damping
Figure 7: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning y$/\kappa_{2}$ for different values of mechanical damping rate $\gamma$. (a) $\gamma/\kappa_{2}=0.001$, (b) $\gamma/\kappa_{2}=0.01$, (c) $\gamma/\kappa_{2}=0.1$, (d) $\gamma/\kappa_{2}=1$. The other parameters are the same as in Figure 2, except for $f/\kappa_{2}=1$, $J_{3}=\sqrt{\frac{-R_{8}+\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$ and $\kappa_{1}/\kappa_{2}=1$.
Figure 8: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning y$/\kappa_{2}$ for different values of mechanical damping rate $\gamma$. (a) $\gamma/\kappa_{2}=0.001$, (b) $\gamma/\kappa_{2}=0.01$, (c) $\gamma/\kappa_{2}=0.1$, (d) $\gamma/\kappa_{2}=1$. The other parameters are the same as in Figure 2 except for $f/\kappa_{2}=1$, $J_{3}=-\sqrt{\frac{-R_{8}-\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$ and $\kappa_{1}/\kappa_{2}=1$.
@@ Acknowledgments
$\mathcal{D}$ $=-2J_{1}|J_{2}|J_{3}G_{2}\cos(\theta-\phi)-2i|J_{2}|J_{3}\kappa_{2}G_{1}\cos\phi$ (31)
@@
$+y^{2}\mathcal{D}_{7}+\kappa_{1}\kappa_{2}\mathcal{D}_{8}-i\gamma fy\mathcal{D}_{9}+y^{4},$
@@
For simplicity, we consider the case where $\theta=\phi=\frac{\pi}{2}$, and the expression of $\mathcal{D}$ becomes $\mathcal{D}=J_{3}^{2}(R_{5}+R_{2}^{\prime})+R_{4}+\frac{R_{6}}{J_{3}^{2}}(J_{1}J_{3}^{2}+J_{1}\gamma f-G_{1}G_{2}f)^{2}+R_{3}$, where $R_{1}=\frac{1}{2\sqrt{\kappa_{1}\kappa_{2}}}$, $R_{2}=\frac{-2G_{1}G_{2}f}{R_{1}}$, $R^{\prime}_{2}=J_{1}^{2}+\kappa_{1}\kappa_{2}$, $R_{3}=G^{2}_{2}f\kappa_{1}+G_{1}^{2}f\kappa_{2}+J_{1}^{2}\gamma f+\kappa_{1}\kappa_{2}\gamma f$, $R_{4}=-2J_{1}f(J_{1}\gamma-G_{1}G_{2})$, $R_{5}=-2J_{1}^{2}$, $R_{6}=\frac{G_{2}^{2}+\gamma\kappa_{2}}{G_{2}^{2}}$, $R_{7}=R_{5}+R_{2}^{\prime}+R_{6}J_{1}^{2}$, $R_{8}=R_{4}-R_{2}+2R_{6}J_{1}^{2}\gamma f-2R_{6}J_{1}G_{1}G_{2}f+R_{3}$, and $R_{9}=R_{6}J_{1}^{2}\gamma^{2}f^{2}-2R_{6}J_{1}\gamma f^{2}G_{1}G_{2}+R_{6}G_{1}^{2}G_{2}^{2}f^{2}$.